Your company has been approached to bid on a contract to sell 3,500 voice recognition (VR) computer keyboards a year for four years. Due to technological improvements, beyond that time they will be outdated and no sales will be possible. The equipment necessary for the production will cost $3.1 million and will be depreciated on a straight-line basis to a zero salvage value. Production will require an investment in net working capital of $88,000 to be returned at the end of the project, and the equipment can be sold for $268,000 at the end of production. Fixed costs are $633,000 per year, and variable costs are $148 per unit. In addition to the contract, you feel your company can sell 8,800, 9,700, 11,800, and 9,100 additional units to companies in other countries over the next four years, respectively, at a price of $275. This price is fixed. The tax rate is 35 percent, and the required return is 9 percent. Additionally, the president of the company will undertake the project only if it has an NPV of $100,000.

  

What bid price should you set for the contract?
Answer #1
0 1 2 3 4
Unit 8800 9700 11800 9100
Investment -$       3,100,000
NWC -$            88,000 $      88,000
Salvage $    268,000
Sales $     3,380,470 $   3,627,970 $ 4,205,470 $ 3,462,970
VC -$    1,820,400 -$   1,953,600 -$ 2,264,400 -$1,864,800
FC -$       633,000 -$      633,000 -$     633,000 -$   633,000
Depreciation -$       775,000 -$      775,000 -$     775,000 -$   775,000
EBT $        152,070 $      266,370 $     533,070 $    190,170
Tax (35%) -$         53,225 -$        93,230 -$     186,575 -$     66,560
Profits $          98,846 $      173,141 $     346,496 $    123,611
Cash Flows -$       3,188,000 $        873,846 $      948,141 $ 1,121,496 $ 1,160,811
NPV $     100,071.73

Sales (year 1) = 8,800 x 275 + 3,500 x bid price

Cost (year 1) = 148 x (8,800 + 3,500) = 1,820,400

FC = 633,000 and depreciation = 3,100,000 / 4 = 775,000

Profits = Sales - Costs - FC - Depreciation - Tax

Cash Flows = Investment + NWC + After-tax Salvage + Profits + Depreciation

NPV = NPV (rate, CF1...CF4) - CF0

Using trial and error, we need to bid price such that NPV = $100,000

we get that at bid price = $274.42